OpenStudy (anonymous):

evaluate the expression arctan0 Please help, I'm really confused...

myininaya (myininaya):

\[\arctan(0)=? \text{ implies } \tan(?)=0 , -\frac{\pi}{2}<?<\frac{\pi}{2}\]

myininaya (myininaya):

So is there any value for ? between -pi/2 and pi/2 such that tan(?)=0

OpenStudy (anonymous):

would that be 0?

myininaya (myininaya):

yes tan(0)=0 so arctan(0)=0 :)

OpenStudy (anonymous):

Thank you!! :)